# RoboHanger: Learning Generalizable Robotic Hanger Insertion for Diverse Garments

### v1

@@ Abstract
For the task of hanging clothes, learning how to insert a hanger into a garment is crucial but has been seldom explored in robotics. In this work, we address the problem of inserting a hanger into various unseen garments that are initially laid out flat on a table. This task is challenging due to its long-horizon nature, the high degrees of freedom of the garments, and the lack of data. To simplify the learning process, we first propose breaking the task into several stages. Then, we formulate each stage as a policy learning problem and propose low-dimensional action parameterization. To overcome the challenge of limited data, we build our own simulator and create 144 synthetic clothing assets to effectively collect high-quality training data. Our approach uses single-view depth images and object masks as input, which mitigates the Sim2Real appearance gap and achieves high generalization capabilities for new garments. Extensive experiments in both simulation and the real world validate our proposed method. By training on various garments in the simulator, our method achieves a 75% success rate with 8 different unseen garments in the real world.
@@ I Introduction
Laundry is a common and routine household chore that is time-consuming. The entire process involves many steps, including collecting clothes, washing, drying, hanging, unfolding, and folding. Therefore, automating laundry tasks with robots is highly appealing. Researchers have explored how to enable robots to fold and unfold clothes from arbitrary initial configurations [1, 2, 3], or to hang rigid or deformable objects on racks or hooks [4, 5, 6, 7]. However, we have identified that one crucial step in the laundry process — inserting a hanger into clothes — has been rarely explored by researchers.
@@
In this work, we investigate how to use a dual-arm robot to insert a hanger into various unseen garments, as illustrated in Figure LABEL:fig:teaser. Our task begins with the target garment laid flat on the table. The goal is to insert the hanger into the garment and hang it up in the final state. End-to-end imitation learning methods face significant challenges in achieving satisfactory results due to (1) task complexity: inserting a hanger into a garment involves intricate, highly coordinated dual-arm actions in a long horizontal manner; (2) data scarcity: collecting large amounts of real-world demonstrations is inefficient and potentially dangerous, while existing simulators lack the necessary physical accuracy and diversity in suitable garments and hangers; and (3) generalizability: garments vary widely in color, shape, size, and initial configurations, making it difficult for models to generalize across different scenarios.
To address the three aforementioned issues, we propose a method that breaks the task into several stages. An overview of the entire process is provided in Figure 1. For different stages, we leverage both learnable and non-learnable modules to take the best of both worlds. Learnable modules offer more flexible and adaptive actions, while non-learnable modules reduce learning difficulty and improve robustness. We design neural networks to learn actions that need to adapt to new clothing shapes and initial configurations, including pressing, lifting, dragging, and rotating, as detailed in Section III-C. For actions that are less critical to the task’s success, we use fixed motions, such as insertion and pulling actions.
We build a simulator and create over 1,000 hanger and 100 garment assets for data collection. Our simulation environment is built upon Taichi [8], which supports large-scale parallel simulations on GPU. We choose depth and the object’s mask as input instead of raw RGB data to mitigate generalization issues. Depth images provide the necessary geometric information for the policies and are naturally unaffected by the varying colors and patterns of different garments. Extensive experiments in the simulation and real world validate our proposed method. Trained on 120 different garments in the simulation, our method can achieve a final success rate of 75% on 8 different unknown garments in the real world without any fine-tuning with real world data.
@@ II Related Work
II-A Deformable Object Manipulation
Deformable object manipulation is a long-standing challenging topic due to the infinite object states and complex dynamics. Some comprehensive surveys can be found in [9, 10, 11]. One type of deformable object manipulation task involves handling one-dimensional objects such as ropes and cables [12, 13, 14, 15, 16, 17]. 2D clothing-like object manipulation tasks include cloth smoothing [18, 19, 20, 1, 21], cloth folding [22, 23, 2], cloth canonicalization [24, 3], bedding manipulation [25, 26] and bag [27, 28, 29] manipulation. Other works study 3D deformable object manipulation, such as solid object [30, 31] manipulation.
From the aspect of the method, early studies often use the model-based method for planning [32, 33, 34]. These kinds of methods often suffer from inaccurate dynamic models. Some methods have also proposed learning the dynamics model or representations of the deformable objects [35, 36, 37, 38, 39]. Recently, more and more research has explored data-driven approaches for learning policies such as imitation learning [40, 41, 42, 43] and reinforcement learning [6, 44, 45, 46]. Some researchers have also used differentiable simulators for deformable object manipulation tasks [47, 48]. In this work, we employ a data-driven approach, utilizing self-supervised trials to train our value network.
II-B Garment Manipulation
Garment manipulation is an important topic in deformable object manipulation. Folding [49, 50, 2] and unfolding [51, 52, 53, 54, 55] are two basic and widely studied directions. Many works also investigated robotic assisted dressing [56, 57, 58, 59]. Hanging garments is an extremely under-explored area. Some related works include: [6] studied using reinforcement learning to hang a fabric on a rack, [7] learned to hang a knit cap on a coat rack, [60] used a rule based method to manipulate the hanger. In this work, we use learning-based method to enable the robot to insert the hanger into the clothes through the neckline.
@@ III Method
We focus on the task of inserting a hanger into a garment. As illustrated in Figure LABEL:fig:teaser, the initial phase involves a hanger and a garment placed separately on the table in front of the robot. The goal is to control a dual-arm mobile robot that inserts the hanger into the garment through the neckline. The final phase involves the robot lifting the hanger with the garment successfully hanging. To focus on the hanger insertion task, we implement simple heuristic policies for the first and final phases, especially, placing the garment faced up to allow its neckline observable. We also omit the control of the robot’s base, as it is beyond the scope of this work. The robot is manually positioned at the table and remains stationary during execution. Each of the robot’s 7-DoF arms is equipped with parallel grippers, and a depth camera is mounted on its head for perception, as shown in Figure III.
Figure 1: System Overview. Our system takes RGB-D observations. First, we segment the hanger, the garment, and its neckline, and detect the three keypoints of the hanger. The depth map, along with the masks, is then fed into our action primitive networks $Q_{1-4}$. Each network predicts a 2D keypoint $P_{1-4}$ in the image, indicating where the action primitives should operate. These keypoints are first back-projected into 3D space and then sent to the whole-body motion controller to execute the long-horizontal hanger insertion task step-by-step.
We will illustrate the overall pipeline in Section III-A and describe each module including: hanger detection in Section III-B, hanger insertion policy in Section III-C, and action primitive networks in Section III-D. The simulation and data collection process is given in Section III-E and the whole body motion controller is introduced in Section III-F.
@@ III-A System Overview
Our method takes RGB-D images as input, as illustrated in Figure 1. We first detect three keypoints of the hanger: a grasping point, and the left and right endpoints. The grasping point indicates where the gripper should grasp the hanger for subsequent motions, while the left and right endpoints mark the locations where the hanger should be inserted into the target garment. Next, we use GroundedSAM [61] to segment the hanger, garment, and its neckline. When the raw depth data is of low quality, we use [62] to obtain a refined depth map. The refined depth map, along with the masks, is then concatenated and fed into action primitive networks to predict two critical action primitives: press-and-lift and drag-and-rotate, which are used to insert the left and right keypoints of the hanger into the garment, respectively.
@@ III-B Garment Segmentation and Hanger Detection
Given a frame of aligned color and depth images, we first use GroundedSAM [61] to segment and obtain the hanger’s mask. Then, using the camera’s intrinsic and extrinsic parameters, we also obtain the hanger’s point cloud as shown in Figure 1. We design a simple network consisting of a PointNet++ backbone and a 3-layer MLP detection head which outputs three keypoints $\{H_{g},H_{l},H_{r}\}$. $H_{g}$ is the grasping point and $H_{l},H_{r}$ denotes the left and right endpoint of the hanger respectively. To train this network, We synthesize 1,000 meshes of hangers with annotated keypoints.
@@ III-C Hanger Insertion Policy
Action primitives [1, 28, 63] are parameterized end-effector motions specifically designed to accomplish robotic tasks. They are effective for learning and robust in execution. Our hanger insertion policy consists of two key action primitives: press-and-lift and drag-and-rotate, each responsible for inserting the left and right ends of the hanger into the garment. Each action primitive is parameterized by two keypoints. Such action primitives are carefully designed to simplify learning and improve the success rate.
@@
As illustrated in Figure 2, after successfully grasping the hanger, we first use the left endpoint $H_{l}$ of the hanger to press the neckline at point $P_{1}$. With the garment pressed by the hanger, we then lift the garment at point $P_{2}$ with the other hand without moving the garment. Finally, we insert the hanger into the garment to a fixed distance. By combining the ’press’ and ’lift’ actions, we can overcome the challenge of the gripper grasping both layers of the garment, which may hinder the insertion of the hanger.
Figure 2: Action primitive: press-and-lift. After grasping the hanger with the right hand, we first press the garment at point $P_{1}$ with the hanger, then lift the garment at point $P_{2}$, and finally insert the hanger to a fixed distance.
@@
In the second action primitive, as illustrated in Figure 3, we first move the left gripper to the position where the right gripper previously released the hanger and re-grasp the hanger to avoid self-collision between the two arms. Next, we use the right gripper to grasp the clothing at point $P_{3}$ and drag it forward to the right. Then, we use the left gripper to rotate the hanger while moving the right end to $P_{4}$ and aligning the angle. Finally, in real-world experiments, we slightly pull the clothing back to arrange it properly.
Figure 3: Action primitive: drag-and-lift. We first switch to the left hand to re-grasp the hanger, avoiding self-collision during subsequent actions. Next, we use the right hand to drag the garment and rotate the hanger into it. Finally, we pull the garment to arrange it properly.
Success criteria and reward
@@
For each action primitive, the success criterion is whether the hanger’s endpoint ($H_{l}$ for the first action primitive and $H_{r}$ for the second) is inserted into the garment. Similarly, if both endpoints are inserted, the task is considered successfully completed. This success criterion also serves as a sparse reward for training our action primitive networks. In simulation, we cast multiple rays randomly from the endpoint in all directions. If the proportion of rays intersecting with the garment’s mesh exceeds a threshold (0.95 for the left endpoint and 0.9 for the right in our experiments), we consider the endpoint to be covered by the garment.
III-D Action Primitive Networks
@@
We use separate value networks to learn the parameters (operation points) for the two action primitives. Specifically, each network learns an action value map using a UNet architecture [64], as illustrated in Figure 1. Each pixel in this map corresponds to the action value of operating on that point. The network input, denoted as $O$, consists of a top-down depth map and object masks. Specifically, we use the garment mask and neckline mask when predicting $P_{1}$ and $P_{2}$ for the press-and-lift action, and the garment mask and hanger mask when predicting $P_{3}$ and $P_{4}$ for the drag-and-rotate action.
To train the four value networks, we simulate over 10,000 trajectories and their corresponding ground truth rewards. Each data entry is of the form $(O,P_{1},P_{2},R)$ where $P_{1}=(X_{1},Y_{1})$ and $P_{2}=(X_{2},Y_{2})$ correspond to each pixel in the output value map, and $R$ is the sparse reward received after executing these actions. The second value network $Q_{2}$ can be directly supervised using the sparse reward $R$ at the end of the execution of both key steps:
$\mathcal{L}_{2}=BCE(Q_{2}[P_{2};O,P_{1}],R)$ (1)
where $BCE$ denotes the binary cross-entropy loss and []delimited-[][$\cdot$][ ⋅ ] is the indexing operation. The input to the $Q_{2}$ is the observation $O$ concatenated with the one-hot map of $P_{1}$. The $Q_{2}$ network serves two purposes here: first, it can predict the value map for $P_{2}$ given $P_{1}$; second, the maximum of its output indicates the action value of point $P_{1}$. Therefore, we propose to use the maximum of the second network’s output $\hat{R}=(\max\{Q_{2}[\cdot;O,P_{1}]\}>0.5)$ as pseudo label for $Q_{1}$. Here $Q_{2}$ is the output after the Sigmoid function. The final loss function for the first network $Q_{1}$ is:
$\mathcal{L}_{1}=BCE(Q_{1}[P_{1};O],\hat{R})$ (2)
We apply a similar data collection and training strategy to train the value networks $Q_{3}$ and $Q_{4}$. When executing these action primitives, we back-project pixels into three-dimensional coordinates using depth value.
III-E Data Collection
We observe that existing simulators, such as PyFlex [65] and MuJoCo [66], either lack the accuracy or efficiency needed for our task. Their main issue is that the self-collisions in the clothing is not well handled. Therefore, we build our simulation environment based on Taichi [8].

### v4

@@ Abstract
For the task of hanging clothes, learning how to insert a hanger into a garment is a crucial step, but has rarely been explored in robotics. In this work, we address the problem of inserting a hanger into various unseen garments that are initially laid flat on a table. This task is challenging due to its long-horizon nature, the high degrees of freedom of the garments and the lack of data. To simplify the learning process, we first propose breaking the task into several subtasks. Then, we formulate each subtask as a policy learning problem and propose a low-dimensional action parameterization. To overcome the challenge of limited data, we build our own simulator and create 144 synthetic clothing assets to effectively collect high-quality training data. Our approach uses single-view depth images and object masks as input, which mitigates the Sim2Real appearance gap and achieves high generalization capabilities for new garments. Extensive experiments in both simulation and reality validate our proposed method. By training on various garments in the simulator, our method achieves a 75% success rate with 8 different unseen garments in the real world.
@@ I Introduction
As a common household task, using a hanger to hang clothes is a challenging deformable manipulation problem for robots. It involves complex contact dynamics between the hanger and the clothes and requires precise coordination between two robotic arms. The entire task consists of flattening the clothes, inserting the hanger, and hanging the hanger on a rack. Researchers have explored methods to flatten clothes from arbitrary initial configurations [1, 2], as well as how to hang objects on racks or hooks [3, 4, 5]. However, we have identified that one crucial step—inserting the hanger into the garment—has rarely been studied.
In recent years, learning-based approaches for deformable object manipulation tasks have made significant progress. Zhao et al. [6] employ end-to-end imitation learning and collect data through teleoperation in the real world to accomplish this task. However, this approach requires substantial human labor for data collection during the training process.
@@
In this work, we investigate how to use a dual-arm robot to insert a hanger into various unseen garments, as illustrated in Figure LABEL:fig:teaser, employing a Sim2Real approach and without any human labeling effort. To focus on the hanger insertion process, the task begins with the target garment laid flat on the table, a scenario that has been addressed in several previous works [1, 2]. This task faces three significant challenges: (1) task complexity: the action space is large and it is difficult to define a good reward function, which makes it hard to find a good policy through sampling by directly applying reinforcement learning algorithms; (2) data scarcity: existing simulators lack the necessary physical accuracy and diversity in terms of suitable garments and hangers; and (3) generalizability: garments vary widely in color, shape, and size, making generalization across different garments a non-trivial problem.
To address the three aforementioned challenges, we propose a method that decomposes the task into two subtasks: inserting the left end of the hanger and inserting the right end of the hanger. Each subtask is accomplished using an action primitive [7, 8, 9], which is a simple trajectory parameterized by two 2D keypoints. These four 2D keypoints are learned through four UNet [10] models. This low-dimensional parameterization facilitates the learning process and ensures robustness during policy execution in both simulation and reality.
To train the models, we build a simulator and generate over 1,000 hangers and 100 garments for data collection. Our simulation environment is built upon Taichi [11], which supports large-scale parallel simulations on GPUs. We generate training data through trial and error within the simulator, minimizing the need for human labeling effort.
We choose depth and the object’s mask as input instead of raw RGB data to mitigate generalization issues. Depth images provide the necessary geometric information for the policies and are naturally unaffected by the varying colors and patterns of different garments.
Extensive experiments in simulation and the real world validate our proposed method. Trained on 120 different garments in simulation, our method can achieve a final success rate of 75% on 8 different unseen garments in reality without any fine-tuning with real data.
@@ II Related Work
II-A Learning-Based Cloth Manipulation
Learning-based robotic cloth manipulation policies have recently gained wide interest. Many works collect data from human demonstrations [6, 9, 12, 1]. The advantage of these methods is the absence of a domain gap, but they require a large amount of manual effort and often suffer from limited generalization. Sim-to-real transfer is a safe and efficient approach [2, 13, 14, 15]. However, previous tasks typically focus only on manipulating the garment itself, without requiring the manipulation of objects other than the garment, such as folding [16, 17, 18] and unfolding [19, 2, 20]. These tasks require a lower diversity of actions, making it easier for the robot to learn.
In contrast, inserting a hanger into clothes is a long-horizon task that requires the manipulation of both hanger and garment: first, the hanger must be grasped, then both endpoints of the hanger must be inserted into the clothing, and finally, the hanger must be lifted. To efficiently generate a large amount of high-quality training data, we design action primitives to reduce the action space. We build our own cloth simulator to automatically generate high-quality training data and labels through trial and error in simulation.
II-B Robotic Hanging
Robotic hanging is a task of significant practical importance, yet it has rarely been studied. Previous works have mainly focused on hanging objects on racks or hooks [5, 21, 22], without using hangers for clothing. In these tasks, typically, only contact between stationary rigid bodies and clothing is involved. However, in our task, the robot needs to simultaneously manipulate both the hanger and the clothing, increasing the complexity of the task.
Researchers [23] have also tried using a rule-based method to manipulate the hanger, but the method they use lacks sufficient generalizability. Additionally, the manipulation policy requires the hanger and clothing to be pre-attached to the robotic arm’s end effector. In our task setup, both the garment and the hanger are initially placed on the table, and the grasping pose is output by our policy. Our manipulation policy for robotic hanger insertion demonstrates strong generalizability across different types of garments.
@@ III Method
Figure 1: System Overview. (a) Before each action primitive, our system takes RGB-D observations as input and segments the hanger, the garment, and its neckline. We pre-detect three keypoints of the hanger at the beginning of the policy. (b) The action primitive press-and-lift inserts the left endpoint of the hanger into the target garment. (c) The action primitive drag-and-rotate inserts the right endpoint of the hanger into the target garment. In (b) and (c), four U-Nets ($Q_{press},Q_{lift},Q_{drag},Q_{rotate}$) all take the depth map and masks as input and output 2D value maps indicating the subtask success rate of applying actions at each pixel. We apply argmax to obtain a final single point.
We focus on the task of inserting a hanger into a garment. As illustrated in Figure LABEL:fig:teaser, the initial state consists of a hanger and a garment placed separately on the table in front of the robot. The goal is to control a dual-arm robot to insert the hanger into the garment through the neckline. The final phase involves the robot lifting the hanger, with the garment successfully hanging. To focus specifically on the task of inserting the hanger, we use the method of previous work [2] to initialize the garment. After inserting the hanger into the clothes, we also omit the step of hanging the hanger on a rack. The robot is manually positioned in front of the table, and its base remains stationary during execution. Each of the 7-DoF arms of the robot is equipped with parallel grippers, and a depth camera is mounted on its head for perception, as shown in Figure II-B.
We will illustrate the overall pipeline in Section III-A and describe each module, including hanger detection in Section III-B, hanger insertion policy in Section III-C, success criteria in Section III-D, action primitive networks in Section III-E, the simulation and data collection process in Section III-F, and the whole-body motion controller in Section III-G.
@@ III-A System Overview
Our method takes RGB-D images as input, as illustrated in Figure 1. We first detect three keypoints of the hanger: the grasping point $H_{g}$, the left endpoint $H_{l}$, and the right endpoint $H_{r}$. The grasping point indicates where the gripper should grasp the hanger for subsequent motions, while the left and right endpoints mark the locations where the hanger should be inserted into the target garment. Next, we use GroundedSAM [24] to segment the hanger, garment, and its neckline. The refined depth map, along with the masks, is then concatenated and fed into action primitive networks to predict two critical action primitives: press-and-lift and drag-and-rotate, which are used to insert the left and right keypoints of the hanger into the garment, respectively.
@@ III-B Garment Segmentation and Hanger Detection
Given a frame of aligned color and depth images, we first use GroundedSAM [24] to segment and obtain the hanger mask. Using the intrinsic and extrinsic parameters of the camera, we can also obtain the point cloud of the hanger as shown in Figure 1(a). We design a simple network consisting of a PointNet++ backbone and a 3-layer MLP detection head that outputs three keypoints $\{H_{g},H_{l},H_{r}\}$. To train this network, we synthesize 1,000 meshes of hangers with annotated keypoints. Generating the data and training takes approximately 1 day to run on a single NVIDIA RTX 3090. After predicting these three keypoints, we grasp the hanger using a calculated grasping pose, with the position centered at the point $H_{g}$ and a rotation of $+45^{\circ}$ around the vector $(H_{l}-H_{r})$, as shown in the bottom left of Figure 1(b).
@@ III-C Hanger Insertion Policy
Action primitives [7, 8, 9, 25] are parameterized end-effector motions specifically designed to accomplish complex robotic tasks. They are effective for learning and robust in execution. Our hanger insertion policy consists of two action primitives: press-and-lift and drag-and-rotate, each responsible for inserting the left and right ends of the hanger into the garment.
@@
This action primitive is parametrized by $(p_{press},p_{lift})=(x_{press},y_{press},x_{lift},y_{lift})\in\mathbb{R}^{4}$. As illustrated in Figure 2, after successfully grasping the hanger, we first use the left endpoint $H_{l}$ of the hanger to press the neckline at point $p_{press}$. With the garment pressed by the hanger, we then lift the garment at point $p_{lift}$ with the other hand without moving the garment. Finally, we insert the hanger into the garment at a fixed distance. By combining the press and lift actions, we can overcome the challenge of the gripper grasping both layers of the garment, which may otherwise hinder the insertion of the hanger.
Figure 2: Action primitive: press-and-lift.
@@
This action primitive is parametrized by $(p_{drag},p_{rotate})=(x_{drag},y_{drag},x_{rotate},y_{rotate})\in\mathbb{R}^{4}$. As illustrated in Figure 3, we first move the left gripper to the position where the right gripper previously released the hanger and re-grasp the hanger to avoid self-collision between the two arms. Next, we use the right gripper to grasp the garment at point $p_{drag}$ and drag it forward to the right. Then, we use the left gripper to rotate the hanger while moving the right end to $p_{rotate}$ and aligning it to the correct angle. Finally, we slightly pull the garment back to arrange it properly.
Figure 3: Action primitive: drag-and-rotate.
III-D Success Criteria
@@
For each action primitive, the success criterion is whether the hanger’s endpoint ($H_{l}$ for the first action primitive and $H_{r}$ for the second) is inserted into the garment. Similarly, if both endpoints are inserted, the task is considered completed. We use $S_{1}$ and $S_{2}$ to represent the success status of each action primitive. $S_{1}=1$ indicates the successful insertion of the left endpoint, and $S_{2}=1$ indicates the successful insertion of the right endpoint. $S_{1}=0$ or $S_{2}=0$ indicates failure.
To detect successful insertion during simulation, we cast multiple rays randomly from the endpoint in different directions. If the proportion of rays intersecting with the garment’s mesh exceeds a threshold (0.95 for the left endpoint and 0.9 for the right in our experiments), we consider the endpoint to be covered by the garment. This proportion is similar to the generalized winding number [26] in $\mathbb{R}^{3}$ space. The thresholds are set manually by inspecting many successful and unsuccessful trajectories and meshes in simulation.
III-E Action Primitive Networks
For each action primitive, we utilize two networks to predict the actions of the left and right hands, respectively. In the first action primitive, we use the network $Q_{press}$ to predict the press position $p_{press}$, and the network $Q_{lift}$ to predict the lift position $p_{lift}$. Likewise, for the second action primitive, we use network $Q_{drag}$ to predict the drag position $p_{drag}$, and network $Q_{rotate}$ to predict the rotate position $p_{rotate}$. These networks are similar to action-value functions in reinforcement learning, but we only consider the success rate within a single subtask.
@@ III-E Action Primitive Networks
The network architecture is based on UNet [10]. The network input, denoted as $O$, consists of a depth image and object masks. Specifically, we use the garment mask and neckline mask when predicting $p_{press}$ and $p_{lift}$ for the press-and-lift action, and the garment mask and hanger mask when predicting $Q_{drag}$ and $Q_{rotate}$ for the drag-and-rotate action. The network output is an image of the same size as the input, with each pixel representing the success rate of inserting one endpoint of the hanger, as mentioned in III-D.
We employ the following dataflow to collaboratively learn two networks in an action primitive. During inference, $Q_{press}$ predicts $p_{press}$, and then $Q_{lift}$ predicts $p_{lift}$ conditioning on $p_{press}$. During training, we use trained $Q_{lift}$ to provide supervision for $Q_{press}$. Specifically, given data of the form $(O_{1},p_{press},p_{lift},S_{1})$, where $O\in\mathbb{R}^{C\times H\times W}$ is the observation, $p_{press},p_{lift}\in\mathbb{R}^{2}$, $S_{1}\in\{0,1\}$ denotes success or failure. We use $Q[y;x]\in\mathbb{R}$ to represent the input image $x$ to the UNet, and select the $y$-th pixel in the output image. The loss function of $Q_{lift}$ is:
$\mathcal{L}_{lift}=BCE(Q_{lift}[p_{lift};O_{1},p_{press}],S_{1})$ (1)
where $BCE$ denotes the binary cross-entropy loss. The input $O_{1}^{\prime}=(O_{1},p_{press})\in\mathbb{R}^{(C+1)\times H\times W}$ to the network $Q_{lift}$ is modified to be $O_{1}$ concatenated with the one-hot map of $p_{press}$, which is used to represent the conditioning on $p_{press}$. The maximum value of the $Q_{lift}$’s output can serve as supervision for $Q_{press}$:
$\hat{S_{1}}=\max\{Q_{lift}[\cdot;O_{1},p_{press}]\}>0.5$ (2)
$\mathcal{L}_{press}=BCE(Q_{press}[p_{press};O_{1}],\hat{S_{1}})$ (3)
Similarly, given data of the form $(O_{2},p_{drag},p_{rotate},S_{2})$, the loss functions of $Q_{drag}$ and $Q_{rotate}$ are:
$\mathcal{L}_{rotate}=BCE(Q_{rotate}[p_{rotate};O_{2},p_{drag}],S_{2})$ (4)
$\hat{S_{2}}=\max\{Q_{rotate}[\cdot;O_{2},p_{drag}]\}>0.5$ (5)
$\mathcal{L}_{drag}=BCE(Q_{drag}[p_{drag};O_{2}],\hat{S_{2}})$ (6)
We apply argmax to the resulting 2D value map to obtain a single final pixel. When executing these action primitives, we back-project pixels into the world frame. The parameters of the action primitives only use the x and y coordinates, while the z coordinate is fixed. This is to avoid failures caused by depth camera inaccuracies.
III-F Data Collection